a 6/53 lottery involves choosing 6 of the numbers from 1 through 53 and a 5/36 lottery involves choosing 5 of the numbers from 1 through 36. the order in which the numbers are chosen does not matter. which lottery is easier to win? explain your answer
Answer by Boreal(15235)

The first is (6/53)(5/52)(4/51)(3/50)(2/49)(1/48)=720/1.65 x 10^10, without rounding until the end, the probability is 4.3559 x 10^(-8)
-----
The second is (5/36)(4/35)(3/34)(2/33)(1/32)=120/45,239,040=2.6526 x 10^(-6). The second is much easier to win (almost 61 times more likely)